OpenStudy (anonymous):

log questions:

OpenStudy (anonymous):

\[\log_{2}3.\log_{3}4 \]

OpenStudy (anonymous):

is equal to (A)1 (B)2 (C)3 (d)4

OpenStudy (anonymous):

pls help

terenzreignz (terenzreignz):

You need two properties... (not the most well-known log properties, too bad)

terenzreignz (terenzreignz):

The first one being \[\Large \log_ab = \frac{1}{\log_ba}\]

OpenStudy (unklerhaukus):

\[\large\boxed{\log_b x=\frac{\log_c x}{ \log_c b}}\]

terenzreignz (terenzreignz):

The second one being what Unkle has already posted :D

terenzreignz (terenzreignz):

I want a box... \[\Large \boxed{ \log_ab = \frac{1}{\log_ba}}\]

terenzreignz (terenzreignz):

Well, it turns out, the so-called property I posted follows directly from what Unkle posted.... meh, go figure :D

OpenStudy (unklerhaukus):

im not sure why you need the one you posted @terenzreignz

terenzreignz (terenzreignz):

I was thinking \[\Large \log_23\cdot\log_34 = \frac{\log_34}{\log_32}=\log_24\]

OpenStudy (anonymous):

\[1/\log_{3}2 . \log_{3}4 \]

OpenStudy (unklerhaukus):

ah, i was thinking \[\log_23\cdot\log_34=\frac{\cancel{\log 3}}{\log2}\cdot\frac{\log 2^2}{\cancel{\log 3}}\]

terenzreignz (terenzreignz):

Many ways to climb the mountain :D Signing off... ^.^

OpenStudy (anonymous):

@UnkleRhaukus which property is that

OpenStudy (anonymous):

and what will be the answer from the options

OpenStudy (unklerhaukus):

pardon?

OpenStudy (precal):

I guess McLove is looking for the final solution

OpenStudy (anonymous):

yes its true i am just verifying my answer

OpenStudy (anonymous):

and thanks

OpenStudy (unklerhaukus):

what did you get @McLove ?

OpenStudy (anonymous):

2

OpenStudy (precal):

you are correct

OpenStudy (anonymous):

thanks guys

OpenStudy (precal):

anytime :)

OpenStudy (agent0smith):

@terenzreignz I've never seen the property you posted, but it makes sense \[\Large \log_ab = \frac{ \log_b b }{ \log_b a } = \frac{1}{\log_ba}\]